OpenStudy (anonymous):

Will give fan and medal! PLEASE help! Match each equation with its y-intercept. 1.) Negative three x plus two y equals twelve. 2.) Three x minus two y equals twelve. 3.)Two x minus three y equals twelve. 4.)Negative two x plus three y equals twelve. A) Negative four B) Negative six C) 6 D) 4

OpenStudy (anonymous):

To get the x-intercept, set y to zero and solve for x. 3x - 2(0) = 12 3x = 12 x = 4 2x - 3(0) = 12 2x = 12 x = 6 -3x + 2(0) = 12 -3x = 12 x = -4 -2x + 3(0) = 12 -2x = 12 x = -6 (2a + 1) + 3b - 6 + a First note that this can be simplified to 3a + 3b - 5 3(3) + 3(7) - 5 = 9 + 21 - 5 = 25 3(-3) + 3(-7) - 5 = -9 - 21 - 5 = -35 3(-3) + 3(7) - 5 = -9 + 21 - 5 = 7 3(3) + 3(-7) - 5 = 9 - 21 - 5 = -17

OpenStudy (aum):

Here you are asked the y-intercept. Set x = 0 and solve for y.

OpenStudy (tkhunny):

Or... 1.) Negative three x plus two y equals twelve. \(-3x + 2y = 12 \longrightarrow \dfrac{x}{-4}+\dfrac{y}{6} = 1\) 2.) Three x minus two y equals twelve. \(3x - 2y = 12 \longrightarrow \dfrac{x}{4}+\dfrac{y}{-6} = 1\) 3.)Two x minus three y equals twelve. \(2x - 3y = 12 \longrightarrow \dfrac{x}{6}+\dfrac{y}{-4} = 1\) 4.)Negative two x plus three y equals twelve. \(-2x + 3y = 12 \longrightarrow \dfrac{x}{-6}+\dfrac{y}{4} = 1\) Look at the denominators associated with y. It is a beautiful thing.